A student receive test scores of 62, 83, and 91. The student's term project score is 88 and her homework score is 76. Each test is worth 20% of the final grade, the term project is 25% of the final grade, and the homework grade is 15% of the final grade. What is the student's mean score in the class?
Answer #1

The student's mean score = 62 * 0.2 + 83 *0.2 + 91 * 0.2 + 88 * 0.25 + 76 * 0.15 = 80.6 (ans)